OpenStudy (anonymous):

-5 / (16 - x^2) divided by 10 / (x - 4)

OpenStudy (anonymous):

To divide by a fraction, multiply by the reciprocal.

OpenStudy (anonymous):

\[ \div \frac ab\to \times \frac ba \]

OpenStudy (anonymous):

Yes, thank you...i worked the problem out and got 1 / 2 (x + 4)... is that correct...i am trying to see if i am getting the hang of these problems:)

OpenStudy (anonymous):

Where did the negative go?

OpenStudy (anonymous):

The fully simplified answer is \[ -\frac 1 {2x+8} \]

OpenStudy (anonymous):

so to fully simplify i need to distribute so no parenthesis are left

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok thank you...don't you need to multiply by -1 in order to rearrange the 16 - x^2

OpenStudy (anonymous):

Oh, I see. Okay that makes sense.

OpenStudy (anonymous):

so the answer is 1 / 2x + 8

OpenStudy (anonymous):

Yeah that is fine.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

would you mind another question?

OpenStudy (anonymous):

Sure